### v2

@@ Appendix D Single-Action Auctions (Section 5) Appendix
$\left\{v^{\leadsto i}\right\}_{i\in B}=\begin{cases}\left(v_{R},\tilde{v}_{-R}\right)&i\in R\\ \left(v_{R},0\right)&\text{otherwise}\end{cases}.$
By the definition of credibility, the auctioneer can report any value to other bidders when the value reported to him is 00 and bidders with value 00 are told the other bidders’ true reports. Therefore, $\left\{v^{\leadsto i}\right\}\in\mathcal{V}^{\psi}(v_{R},0)$ and
$\sum_{i}\tilde{t}_{i}(v^{\leadsto i})=\sum_{i\in R}\tilde{t}_{i}(v_{R},\tilde{v}_{-R})+\sum_{i\notin R}\tilde{t}_{i}(v_{R},0)>\sum_{i}\tilde{t}_{i}((v_{R},0)).$
This contradicts Lemma O.9, and so the auction must be weakly shill-proof.
$(\psi=\emptyset)$-Credibility $\to$ Strong Shill-Proofness. Suppose that the auction is not strongly shill-proof and $\psi=\emptyset$. There are two cases to consider: either the auction is not weakly shill-proof or it is. If the auction is not weakly shill-proof, then we can apply the previous case to conclude the auction is not $(\psi=\emptyset)$-credible. If the auction is weakly shill-proof, but not strongly shill-proof, then by Lemma O.10, there exists $R,k^{*}\in R,v_{R}$, and $v_{-R}\text{ such that }\tilde{t}_{k^{*}}(v)>\tilde{t}_{k^{*}}(v_{R},0)$. Thus, we can construct the following profitable auctioneer reporting deviation:
$\left\{v^{\leadsto i}\right\}_{i\in B}=\begin{cases}\left(v_{R},v_{-R}\right)&i=k^{*}\\ \left(v_{R},0\right)&\text{otherwise}\end{cases}.$
Since $\psi=\emptyset$, we know that $\left\{\tilde{v}^{\leadsto i}\right\}\in\mathcal{V}^{\psi}(v_{R},0)$. The total transfers is then
$\sum_{i}\tilde{t}_{i}\left(v^{\leadsto i}\right)=\tilde{t}_{k^{*}}(v)+\sum_{i\neq k^{*}}\tilde{t}_{i}\left(v_{R},0\right)>\sum_{i}\tilde{t}_{i}\left(v_{R},0\right);$
hence, by Lemma O.9, we see that the auction is not credible.
Strong Shill-Proofness $\to$ $\psi$-Credibility. Suppose that the auction is not $\psi$-credible. Then, combining Lemma O.9 with the ex-post IR condition, there exists $v,\left\{v^{\leadsto i}\right\}\in\mathcal{V}^{\psi}(v)$ and $k^{*}$ such that $\tilde{t}_{k^{*}}(v^{\leadsto k^{*}})>\tilde{t}_{k^{*}}(v)$. Recall, by the definition of $\psi^{-1}$, that $v^{\leadsto k^{*}}_{k^{*}}=v_{k^{*}}$. Suppose $R=\left\{k^{*}\right\}$. Then,
$\sum_{i\in R}\tilde{t}_{i}(v^{\leadsto k^{*}})=\tilde{t}_{k^{*}}(v^{\leadsto k^{*}})>\tilde{t}_{k^{*}}(v)\geq\tilde{t}_{k^{*}}(v_{k^{*}},0)=\sum_{i\in R}\tilde{t}_{k^{*}}(v_{k^{*}},0).$
Therefore, by Lemma D.3, the auction is not strongly shill-proof. ∎

### v3

@@ E.2 Generalizing Credibility in the Single-Action Case
$\left\{\theta^{\leadsto i}\right\}_{i\in B}=\begin{cases}\left(\theta_{R},\tilde{\theta}_{-R}\right)&i\in R\\ \left(\theta_{R},0\right)&\text{otherwise}\end{cases}.$
By the definition of credibility, the auctioneer can report any value to other bidders when the value reported to him is 0 and bidders with value 0 are told the other bidders’ true reports. Therefore, $\left\{\theta^{\leadsto i}\right\}\in\vartheta^{\psi}(\theta_{R},0)$ and
$\sum_{i}\tilde{t}_{i}(\theta^{\leadsto i})=\sum_{i\in R}\tilde{t}_{i}(\theta_{R},\tilde{\theta}_{-R})+\sum_{i\notin R}\tilde{t}_{i}(\theta_{R},0)>\sum_{i}\tilde{t}_{i}((\theta_{R},0)).$
This contradicts Lemma E.11, and so the auction must be weakly shill-proof.
$(\psi=\emptyset)$-Credibility $\to$ Strong Shill-Proofness. Suppose that the auction is not strongly shill-proof and $\psi=\emptyset$. There are two cases to consider: either the auction is not weakly shill-proof or it is. If the auction is not weakly shill-proof, then we can apply the previous case to conclude the auction is not $(\psi=\emptyset)$-credible. If the auction is weakly shill-proof, but not strongly shill-proof, then by Lemma E.12, there exists $R,k^{*}\in R$, and $\theta\text{ such that }\tilde{t}_{k^{*}}(\theta)>\tilde{t}_{k^{*}}(\theta_{R},0)$. Thus, we can construct the following profitable auctioneer reporting deviation:
$\left\{\theta^{\leadsto i}\right\}_{i\in B}=\begin{cases}\left(\theta_{R},\theta_{-R}\right)&i=k^{*}\\ \left(\theta_{R},0\right)&\text{otherwise}\end{cases}.$
Since $\psi=\emptyset$, we know that $\left\{\tilde{\theta}^{\leadsto i}\right\}\in\vartheta^{\psi}(\theta_{R},0)$. The total transfers are then
$\sum_{i}\tilde{t}_{i}\left(\theta^{\leadsto i}\right)=\tilde{t}_{k^{*}}(\theta)+\sum_{i\neq k^{*}}\tilde{t}_{i}\left(\theta_{R},0\right)>\sum_{i}\tilde{t}_{i}\left(\theta_{R},0\right).$
Hence, by Lemma E.11, we see that the auction is not credible.
Strong Shill-Proofness $\to$ $\psi$-Credibility. Suppose that the auction is not $\psi$-credible. Then, combining Lemma E.11 with the ex-post IR condition, there exists $\theta,\left\{\theta^{\leadsto i}\right\}\in\vartheta^{\psi}(\theta)$ and $k^{*}$ such that $\tilde{t}_{k^{*}}(\theta^{\leadsto k^{*}})>\tilde{t}_{k^{*}}(\theta)$. Recall, by the definition of $\psi^{-1}$, that $\theta^{\leadsto k^{*}}_{k^{*}}=\theta_{k^{*}}$. Suppose $R=\left\{k^{*}\right\}$. Then,
$\sum_{i\in R}\tilde{t}_{i}(\theta^{\leadsto k^{*}})=\tilde{t}_{k^{*}}(\theta^{\leadsto k^{*}})>\tilde{t}_{k^{*}}(\theta)\geq\tilde{t}_{k^{*}}(\theta_{k^{*}},0)=\sum_{i\in R}\tilde{t}_{k^{*}}(\theta_{k^{*}},0).$
Therefore, by Lemma E.13, the auction is not strongly shill-proof. ∎